# Faster Language Models with Better Multi-Token Prediction Using Tensor Decomposition

### v1

@@ Abstract
We propose a new model for multi-token prediction in transformers, aiming to enhance sampling efficiency without compromising accuracy. Motivated by recent work that predicts the probabilities of subsequent tokens using multiple heads, we connect this approach to rank-$1$ canonical tensor decomposition. By generalizing it to a rank-$r$ canonical probability decomposition, we develop an improved model that predicts multiple tokens simultaneously. This model can also be interpreted as a mixture of experts, allowing us to leverage successful techniques from that domain for efficient and robust training. Importantly, the overall overhead for training and sampling remains low. Our method demonstrates significant improvements in inference speed for both text and code generation tasks, proving particularly beneficial within the self-speculative decoding paradigm. It maintains its effectiveness across various model sizes and training epochs, highlighting its robustness and scalability.
@@ 1 Introduction
where $w_{\alpha}\geq 0$ are learnable weights satisfying $\sum_{\alpha=1}^{r}w_{\alpha}=1$.
The proposed formulation in equation 2 accounts for dependencies among future tokens by effectively considering a mixture of expert models, each capturing different aspects of the token distribution. By leveraging this rank-$r$ decomposition, we aim to enhance the accuracy of multi-token predictions, thereby increasing token acceptance rates during speculative decoding and reducing overall inference time. Thus, our main contributions are as follows:
@@ 2.1 Overall concept
$A\in\mathbb{R}^{V\times V\times\ldots\times V},\quad A[x_{t+1},\ldots,x_{t+n}]=P_{\theta}(x_{t+n:t+1}|x_{t:1}),$ (3)
where $V$ is the vocabulary size. The tensor $A$ encapsulates the probabilities of all possible combinations of the next $n$ tokens. In Gloeckle et al. (2024) it was proposed to approximate this joint distribution by assuming that future tokens are conditionally independent given the past as shown in equation 1. We draw special attention to the fact that this may be interpreted as a rank-$1$ CP approximation to the tensor $A$. While computationally efficient, such approximation ignores dependencies among the future tokens.
To better capture these dependencies, we propose to approximate the joint distribution using a rank-$r$ CP tensor decomposition according to equation 2. In order to ensure that $P_{\theta}$ from this equation is indeed a probability tensor, it is sufficient to undertake that
$w_{\alpha}\geq 0,\quad\sum_{\alpha=1}^{r}w_{\alpha}=1.$ (4)
The difference between equation 2 and standard CP-decomposition is an additional constraint on the factors of decomposition, i.e., each factor, $P_{\theta}^{(s)}$ should be non-negative and sum up to $1$ along one mode:
@@ 2.2 Training procedure
In training, we maximize the log-likelihood of the predicted $n$ tokens. The computation of the log-likelihood is straightforward: first, the embeddings are calculated by the transformer backbone (it has the same cost as for the next token prediction). We need to evaluate the logarithm of the likelihood, so using equation 6 directly is not numerically stable. Instead, we compute everything using the logarithms of the probabilities. For each pair of sequences $x_{t:1}$ and $x_{t+n:t+1}$, we evaluate the logarithm of the mixture weights $w$ (the computational cost corresponds to a matrix-by-matrix product and logsoftmax operation), then use equation 6 to compute $n$ matrices of the size $V\times r$
@@
$\begin{split}L&=\log\left(P_{\theta}(x_{t+n:t+1}|x_{t:1})\right)\approx\log\left(\sum_{\alpha=1}^{r}w_{\alpha}\prod_{s=1}^{n}P_{\theta}^{(s)}(x_{t+s}|x_{t:1},\alpha)\right)=\\ &=\log\left(\sum_{\alpha=1}^{r}w_{\alpha}\prod_{s=1}^{n}\exp(C^{(s)}_{\theta,\alpha})\right)=\log\left(\sum_{\alpha=1}^{r}\exp\left(\log{w_{\alpha}}+\sum_{s=1}^{n}C^{(s)}_{\theta,\alpha}\right)\right).\end{split}$ (9)
@@ 2.3 Auxilary load balancing loss
Each term of the summation in equation 9 corresponds to a single expert, which predicts its own probabilities for each token. We have found, that optimizing such loss directly leads to the effects, similar to the ones observed in Mixture Of Experts (MoE) framework (Masoudnia & Ebrahimpour, 2014; Cai et al., 2024): one expert (i.e., rank-1 term in our case) dominates the others, leading to worser likelihood even in the presence of larger number of parameters. Note, that such interpretation and connection is not well-known in the low-rank approximation community, and can be investigated further on. To obtain the balance between different experts, we utilize the achievements from the MoE communities and propose to use an auxiliary balancing loss on $w$.
@@ 2.4 Sampling method
Our sampling scheme is similar to the one proposed in Gloeckle et al. (2024). We sample candidates from the proposal distribution (our approximation to the joint distribution of the next tokens) and then accept them or reject according to the recommendations of the draft model (which is the same model that predicts the next token).
@@
$P_{\theta}(x_{t+2}|x_{t+1})=\sum_{\alpha=1}^{r}w_{\alpha}P_{\theta}^{(1)}(x_{t+1}|x_{t:1},\alpha)P_{\theta}^{(2)}(x_{t+2}|x_{t:1},\alpha),$ (12)
@@
$P_{\theta}(x_{t+s}|x_{t+1},\ldots x_{t+s-1})=\sum_{\alpha=1}^{r}w_{\alpha}\prod_{k=1}^{s-1}P_{\theta}^{(k)}(x_{t+k}|x_{t:1},\alpha).$ (13)
@@ 2.5 Speculative Decoding
Speculative decoding (Chen et al., 2023; Leviathan et al., 2023) is a technique designed to accelerate the inference process of autoregressive models by generating multiple tokens in parallel, thereby reducing the latency associated with sequential token generation. In traditional autoregressive sampling, tokens are generated one at a time, with each new token conditioning on the previously generated tokens. This sequential nature inherently limits the speed of generation, especially for lengthy outputs.
Our sampling method seamlessly integrates with the speculative decoding framework by enhancing its capacity to handle multi-token predictions, as can be seen from Algorithm 1. The usual setup for a speculative decoding consists of a draft model and a base model. In our case we implement a modification of a self-speculative decoding algorithm, as described in Zhang et al. (2023); Elhoushi et al. (2024). So, as a base model we take a next-token prediction model and as a draft model – the prediction for $dim$ tokens forward obtained from a full “CP-head”.
@@
Algorithm 1 Self-speculative decoding with rank-r experts
0:  prefix $X$, encoder $E$, weight function $W$, heads $H_{i}$, dim $n$, rank $r$
1:  $e_{t}\leftarrow E(X)[-1]$Take last embedding}
2:  $w_{t}\leftarrow log(\text{softmax}(W(e_{t})))$Obtain expert weights}
3:  $LP_{t}\leftarrow[H_{i}(e_{t}):0\leq i\leq r]$Obtain log core probabilities by expert}
4:  $S_{t},P_{t}\leftarrow RankRSample(w_{t},LP_{t})$Sample from rank $r$ head as described in equation 2. $S_{t}$ are the samples and $P_{t}$ are the conditional probability distributions}
5:  $e_{p}^{\text{list}}\leftarrow E(X)[-\text{n}:]$Obtain new embeddings}
6:  $\text{accept}\leftarrow True$
7:  $i\leftarrow 0$We use the scheme as in Leviathan et al. (2023) below}
8:  while accept do
9:     $u\leftarrow Uniform(0,1)$
10:     $P^{i}_{p}\leftarrow FirstHeadPrediction(e_{p}^{i},W,H_{1})$Obtain probability distribution from first head as described in equation 2}
11:     $c\leftarrow P^{i}_{p}[S_{t}^{i+1}]/P_{t}^{i+1}[S_{t}^{i+1}]$Different indexes due to offset induced by the fact, that first token by our draft model is always from the same distribution, as it would be from a base model}
12:     if $u<c$ then
13:        $i\leftarrow i+1$
14:     else
15:        $\text{accept}\leftarrow False$
16:     end if
17:  end while
18:  if $i<n$ then
19:     $P_{last}\leftarrow\text{normalize}(max(0,P^{i}_{p}-P_{t}^{i+1}))$
20:  else
21:     $P_{last}\leftarrow FirstHeadPrediction(e_{p}^{n},W,H_{1})$
22:  end if
23:  $s_{l}\leftarrow\text{sample}(P_{last})$additional sample from base model}
24:
25:  return  $\text{prefix}+S_{t}[:i+1]+s_{l}$
@@ 3.1 Training different ranks and auxiliary loss models
For our experiments we’ve chosen the multi-head tiny transformer model with 56.3 M parameters based on the code in Karpathy (2022). We consider the case of $4$ heads and added RoPe positional encodings as in Su et al. (2024). Training was conducted on the Tiny Stories dataset (Eldan & Li, 2023) using various ranks for the CP-decomposition. The objective was to observe how increasing the rank influences the joint loss and loss on the first token. Because the quality of our final generation depends only on the quality of the first head, we tracked both those metrics. Additionally, we experimented with different sizes of the auxiliary loss penalty to ensure balanced expert utilization.
@@
Contrary to the joint loss, right graph in Figure 3 shows that the loss for the first token remains largely unchanged across different ranks. It is worth noting, that probability distribution for the first token as a function of last layer embeddings $m$ is given by $\sum_{\alpha=1}^{r}w_{\alpha}(m)C_{\alpha}(m)$ (in notation of equation 9) and both $C_{\alpha}$ and $w_{\alpha}$ are linear, which makes this function equivalent to a simple linear head. So, after convergence we expected the same loss for all of the ranks. As follows from the reported results, this is exactly what happened and this consistency confirms that our model maintains optimal training for the initial token prediction, ensuring that the foundational aspects of the sequence generation remain robust. The loss on the first token is especially crucial, because with a speculative decoding we are improving sample for a big model, which is in our self-speculative case is a first head. We also note that from Figure 2 it follows that all of the inference speedup will be obtained without compromising quality.
@@
3.2 Head-only fine-tuning for PyCode model
To evaluate the flexibility of our approach, we fine-tuned only the prediction head of the PyCodeGPT (Zan et al., 2022) model across different ranks on the Github Code dataset by CodeParrot. 111 See https://huggingface.co/datasets/codeparrot/github-code. This experiment aimed to determine whether partial model updates could yield performance improvements without the computational overhead of full model fine-tuning.
Figure 5 and Table 2 demonstrates speculative decoding performance for the experiments we conducted for different rank values. From the reported results it follows that even when only the head is fine-tuned, increasing the rank leads to marginal improvements in joint loss. Additionally, we can see that speculative decoding benefits from higher ranks, albeit to a lesser extent (approximately 9% increase in accepted drafts) compared to the full model training.

### v2

@@ Abstract
We propose a new model for multi-token prediction in transformers, aiming to enhance sampling efficiency without compromising accuracy. Motivated by recent work that predicts the probabilities of subsequent tokens using multiple heads, we connect this approach to rank-$1$ canonical tensor decomposition. By generalizing it to a rank-$r$ canonical probability decomposition, we develop an improved model that predicts multiple tokens simultaneously. This model can also be interpreted as a mixture of experts, allowing us to leverage successful techniques from that domain for efficient and robust training. Importantly, the overall overhead for training and sampling remains low. Our method demonstrates significant improvements in inference speed for both text and code generation tasks, proving particularly beneficial within the self-speculative decoding paradigm. It maintains its effectiveness across various model sizes and training paradigms, highlighting its robustness and scalability.
LLM, Low rank decomposition
@@ 1 Introduction
where learnable weights $w_{\alpha}$ ($\alpha=1,2,\ldots,r$) are satisfying
$w_{\alpha}\geq 0,\quad\sum_{\alpha=1}^{r}w_{\alpha}=1.$ (3)
The proposed formulation in (2) accounts for dependencies among future tokens by effectively considering a mixture of expert models, each capturing different aspects of the token distribution. By leveraging this rank-$r$ decomposition, we aim to enhance the accuracy of multi-token predictions, thereby increasing token acceptance rates during speculative decoding and reducing overall inference time. Thus, our main contributions are as follows:
@@
2 Preliminaries
In this section, we introduce the background and the formulation of the task we aim to solve through the approximation of a probability tensor using CP-decomposition. For simplicity, this section assumes a greedy decoding approach; however, all presented methods are generalized to support probabilistic decoding.
Due to the autoregressive nature of sequence generation in large language models, the cost of generating a single new sample is approximately the same as the cost of generating multiple new samples that share a common prefix. This enables the default speculative decoding paradigm  (Chen et al., 2023; Leviathan et al., 2023), which employs a Draft and then Verify mechanism to allow multiple tokens to be generated in one step, realizing lossless acceleration.
Given a model $M$, prefix $(x_{1},x_{2},\ldots,x_{t})$ and a smaller draft model $M_{d}$ alligned with an original model, we can sequentially generate samples $(x^{d}_{t+1},x^{d}_{t+2},\ldots,x^{d}_{t+n})$ from $M_{d}$. Then, we verify these samples by performing a single forward pass through the original model, generating the sequence $(x^{c}_{t+1},x^{c}_{t+2},\ldots,x^{c}_{t+n})$, where $x^{c}_{t+k}=M(x_{1},x_{2},\ldots,x_{t},x^{d}_{t+1},\ldots,x^{d}_{t_{k}-1})$. The longest common prefix between $x^{d}_{t+1:t+n}$ and $x^{c}_{t+1:t+n}$ of the length $k$ becomes $x_{t+1},\ldots,x_{t+k}$. It is easy to see that these tokens are identical to the ones that would be generated by $M$ under the regular sampling procedure. Additionally, the better the alignment between the smaller model $M_{d}$ and the original model $M$, the longer the common prefix, and thus the greater the speedup.
One possible extension of the described standard speculative decoding paradigm to increase the length of the accepted prefix is tree verification, as presented in (Wang et al., 2024; Xiong et al., 2024). In this approach, multiple sequences are generated from the draft model, and then all tokens are verified simultaneously using a tree-like custom attention mask. The structure of the tree can either be static (with a fixed shape) or dynamic.
The draft model $M_{d}$ can also be obtained through modifications of the original model (self-speculative decoding) as in (Zhang et al., 2023; Elhoushi et al., 2024; Hooper et al., 2023). This bypasses the need to train a smaller model from scratch and opens up new possibilities for aligning base and draft model.
One way to obtain a draft model for a self-speculative decoding scenario is to modify the original model so that it predicts multiple tokens at once during training. This approach was implemented in a simple form in (Gloeckle et al., 2024), and more recent works, like Deepseek v3 (Liu et al., 2024), use a more sophisticated method for predicting multiple tokens. From the perspective of the speculative decoding framework, however, the concept remains the same: we have a draft model $M_{d}$, which generates multiple tokens at once, and a target model $M$, where only the parameters responsible for generating the first token are activated.
In all of those cases we want our draft model $M_{d}$ to generate multiple tokens faster than the original model $M$, while ensuring that the generated tokens are as close as possible to the original ones. Therefore, approximating the joint probability $P(x_{t+n:t+1}|x_{t:1})$ becomes an important challenge.
@@ 3.1 Overall concept
$\begin{split}A\in&\mathbb{R}^{V\times V\times\ldots\times V},\\ A[x_{t+1},\ldots,x_{t+n}]=&P_{\theta}(x_{t+n:t+1}\,|\,x_{t:1}),\end{split}$ (4)
where $V$ is the vocabulary size. The tensor $A$ encapsulates the probabilities of all possible combinations of the next $n$ tokens. In (Gloeckle et al., 2024) it was proposed to approximate this joint distribution by assuming that future tokens are conditionally independent given the past as shown in (1). We draw special attention to the fact that this may be interpreted as a rank-$1$ CP approximation to the tensor $A$. While computationally efficient, such approximation ignores dependencies among the future tokens.
To better capture these dependencies, we propose to approximate the joint distribution using a rank-$r$ CP tensor decomposition according to (2). In order to ensure that $P_{\theta}$ from this equation is indeed a probability tensor, it is sufficient to fulfill the normalization condition (3). The difference between (2) and standard CP-decomposition is an additional constraint on the factors of decomposition, i.e., each factor, $P_{\theta}^{(s)}$ should be non-negative and sum up to $1$ along one mode:
@@ 3.2 Training procedure
In training, we maximize the log-likelihood of the predicted $n$ tokens. The computation of the log-likelihood is straightforward: first, the embeddings are calculated by the transformer backbone (it has the same cost as for the next token prediction). We need to evaluate the logarithm of the likelihood, so using (6) directly is not numerically stable. Instead, we compute everything using the logarithms of the probabilities. For each pair of sequences $x_{t:1}$ and $x_{t+n:t+1}$, we evaluate the logarithm of the mixture weights $w$ (the computational cost corresponds to a matrix-by-matrix product and logsoftmax operation), then use (6) to compute $n$ matrices of the size $V\times r$
@@
$\begin{split}L&=\log\left(P_{\theta}(x_{t+n:t+1}|x_{t:1})\right)\\ &\approx\log\left(\sum_{\alpha=1}^{r}w_{\alpha}\prod_{s=1}^{n}P_{\theta}^{(s)}(x_{t+s}|x_{t:1},\alpha)\right)\\ &=\log\left(\sum_{\alpha=1}^{r}w_{\alpha}\prod_{s=1}^{n}\exp(C^{(s)}_{\theta,\alpha})\right)\\ &=\log\left(\sum_{\alpha=1}^{r}\exp\left(\log{w_{\alpha}}+\sum_{s=1}^{n}C^{(s)}_{\theta,\alpha}\right)\right).\end{split}$ (9)
@@ 3.3 Auxilary load balancing loss
Each term of the summation in (9) corresponds to a single expert, which predicts its own probabilities for each token. We have found, that optimizing such loss directly leads to the effects, similar to the ones observed in Mixture Of Experts (MoE) framework (Masoudnia & Ebrahimpour, 2014; Cai et al., 2024): one expert (i.e., rank-1 term in our case) dominates the others, leading to worser likelihood even in the presence of larger number of parameters.111 We note, that such interpretation and connection is not well-known in the low-rank approximation community, and can be investigated further on. To obtain the balance between different experts, we utilize the achievements from the MoE communities and propose to use an auxiliary balancing loss on $w$.
@@ 3.4 Sampling method
Our sampling scheme is similar to the one proposed in (Gloeckle et al., 2024). We sample candidates from the proposal distribution (our approximation to the joint distribution of the next tokens) and then accept them or reject according to the recommendations of the draft model (which is the same model that predicts the next token).
@@
$\begin{split}P_{\theta}(x_{t+2}&\,|\,x_{t+1})=\\ \sum_{\alpha=1}^{r}&w_{\alpha}P_{\theta}^{(1)}(x_{t+1}|x_{t:1},\alpha)P_{\theta}^{(2)}(x_{t+2}|x_{t:1},\alpha),\end{split}$ (12)
@@
$\begin{split}P_{\theta}(x_{t+s}\,|\,&x_{t+1},\ldots x_{t+s-1})=\\ &\sum_{\alpha=1}^{r}w_{\alpha}\prod_{k=1}^{s-1}P_{\theta}^{(k)}(x_{t+k}|x_{t:1},\alpha).\end{split}$ (13)
@@ 3.5 Speculative decoding
One of the advantages of multi-token models is the ability to accelerate the decoding of the base model in speculative decoding paradigm without compromising accuracy (lossless acceleration). Our sampling method seamlessly integrates with the speculative decoding framework by enhancing its capacity to handle multi-token predictions. The usual setup for a speculative decoding consists of a draft model and a base model. In our case we implement a modification of a self-speculative decoding algorithm, as described in (Zhang et al., 2023; Elhoushi et al., 2024). So, as a base model we take a next-token prediction model and as a draft model – the prediction for $dim$ tokens forward obtained from a full “CP-head”. In some experiments we also use a simple static draft tree, where for each token we sample top-$n$ candidates. Any advanced tree building strategy such as (Miao et al., 2024; Xiong et al., 2024) can also be used.222 Our work focuses on “better draft model” part of speculative decoding, not on the “most effective token tree” part, so we keep the token generation part simple.
@@
3.6 Fine tuning
In the case, when we already have a pretrained model $M$ and want to produce $M_{d}$ out of it by training only the new tensor head, we face a different challenges, as in the case, when entire model is trained from scratch. To address those challenges we propose two main modifications.
First one is the design of our custom tensor head. When we already have an initial head $W\in\mathbb{R}^{V\times E}$ then instead of taking $W^{(s)}_{\alpha}\in\mathbb{R}^{V\times E}$, we can use a shared head matrix $W$ a smaller $w^{(s)}_{\alpha}\in\mathbb{R}^{V\times E}$, and set $W^{(s)}_{\alpha}=W\times w^{(s)}_{\alpha}$. The difference can be seen in Figure 2. While this theoretically reduces the expressivity of our custom head, this approach is highly efficient, as it’ll be demonstrated in the next section. Also, since the embedding size $E$ is usually at least an order of magnitude lower, than vocabulary size $V$, the total amount of parameters in tensor head is greatly reduced.
The second one is the change in the training paradigm. Since now one have an access to the logits of the target model and those logits are unchanged during the training, one can use distillation loss in combination with targets given by an inital sequence. Using the fact, that we want distribution, generated by our head be as close to initial head as possible, we write the distribution for the $k$-th token like
$p^{d}_{t+k}\thicksim P_{\theta}^{k}(x_{t+k}|x_{1:t},p^{c}_{t+1:t_{k}-1}),$
where $p^{c}_{t+k}\thicksim P(x_{t+k}|x_{1:t+k-1})$ (distribution given by original head with respect to the previous tokens). Loss for the $k$-th token will look like
$L_{k}=\beta\cdot KL(p^{d}_{t+k},p^{c}_{t+k})+(1-\beta)\cdot CE(p^{d}_{t+k},x_{t+k}),$ (14)
where $\beta$ is a constant set to 0.9, $KL$ is a Kullback–Leibler divergence and $CE$ is a Cross-entropy. Then, the full loss looks like
$L=\sum_{i=0}^{d}\left(\gamma^{i}L_{i}\right)+\mathcal{L}_{\text{aux}},$ where $\gamma$ is a discount parameter and $\mathcal{L}_{\text{aux}}$ is the same, as in case of training from scratch. This loss is inspired by the usual distillation logic and the loss used by (Gao et al., 2024).
Figure 2: Comparison of original (left) and reduced (right) tensor head designs. $W$ is unchanged during fine-tuning process.
@@ 4.1 Training different ranks and auxiliary loss models
For our experiments we’ve chosen the multi-head tiny transformer model with 56.3 M parameters based on the code in (Karpathy, 2022). We consider the case of $4$ heads and added RoPe positional encodings as in (Su et al., 2024). Training was conducted on the Tiny Stories dataset (Eldan & Li, 2023) using various ranks for the CP-decomposition. The objective was to observe how increasing the rank influences the joint loss and loss on the first token. Because the quality of our final generation depends only on the quality of the first head, we tracked both those metrics. Additionally, we experimented with different sizes of the auxiliary loss penalty to ensure balanced expert utilization.
@@
Contrary to the joint loss, right graph in Figure 4 shows that the loss for the first token remains largely unchanged across different ranks. It is worth noting, that probability distribution for the first token as a function of last layer embeddings $m$ is given by $\sum_{\alpha=1}^{r}w_{\alpha}(m)C_{\alpha}(m)$ (in notation of (9)) and both $C_{\alpha}$ and $w_{\alpha}$ are linear, which makes this function equivalent to a simple linear head. So, after convergence we expected the same loss for all of the ranks. As follows from the reported results, this is exactly what happened and this consistency confirms that our model maintains optimal training for the initial token prediction, ensuring that the foundational aspects of the sequence generation remain robust. The loss on the first token is especially crucial, because with a speculative decoding we are improving sample for a big model, which is in our self-speculative case is a first head. We also note that from Figure 3 it follows that all of the inference speedup will be obtained without compromising quality.
@@
To ensure scalability of our approach we’ve also conducted experiments on a bigger model. Namely, we took an 1B model, obtained by increasing the amount of hidden size and amount of transformer blocks of the model, implemented in (GPT-2 architecture) and tuned the obtained model on the fineweb ((Penedo et al., 2024b)) dataset. The results are presented in the Table 1. We’ve achieved 24% increase in accepted draft length, when compared to a rank-1 model and 44% reduction in execution time, when compared to a baseline.
Table 1: Inference time comparison for 1B GPT2-like model
Model Time (s) Param. (B) Avg. acceptance
base 43.8 - -
rank 1 34.0 0.98 1.58
rank 2 28.9 1.2 1.96
rank 4 31.5 1.9 1.85
@@
4.2 Fine-tuning of pretrained model
To evaluate the flexibility of our approach, we fine-tuned only the prediction head of the 1.82B model provided by the (Penedo et al., 2024a) using the $fineweb$-$edu$ dataset. This experiment aimed to determine whether partial model updates could yield performance improvements without the computational overhead of full model fine-tuning.
Figure 6 and Table 3 demonstrates speculative decoding performance for the experiments we conducted for different rank values. The results were obtained on temperature 0 using a simple uniform draft tree with 5 branches. From the reported results it follows that even when only the head is fine-tuned, increasing the rank leads to marginal improvements in joint loss. Additionally, we can see that speculative decoding benefits from higher ranks, albeit to a lesser extent (approximately 13% increase in accepted drafts) compared to the full model training.